# HypeRL: Parameter-Informed Reinforcement Learning for Parametric PDEs

### v1

@@ Abstract.
In this work, we devise a new, general-purpose reinforcement learning strategy for the optimal control of parametric partial differential equations (PDEs). Such problems frequently arise in applied sciences and engineering and entail a significant complexity when control and/or state variables are distributed in high-dimensional space or depend on varying parameters. Traditional numerical methods, relying on either iterative minimization algorithms – exploiting, e.g., the solution of the adjoint problem – or dynamic programming – also involving the solution of the Hamilton-Jacobi-Bellman (HJB) equation – while reliable, often become computationally infeasible. Indeed, in either way, the optimal control problem has to be solved for each instance of the parameters, and this is out of reach when dealing with high-dimensional time-dependent and parametric PDEs. In this paper, we propose HypeRL, a deep reinforcement learning (DRL) framework to overcome the limitations shown by traditional methods. HypeRL aims at approximating the optimal control policy directly, bypassing the need to numerically solve the HJB equation explicitly for all possible states and parameters, or solving an adjoint problem within an iterative optimization loop for each parameter instance. Specifically, we employ an actor-critic DRL approach to learn an optimal feedback control strategy that can generalize across the range of variation of the parameters. To effectively learn such optimal control laws for different instances of the parameters, encoding the parameter information into the DRL policy and value function neural networks (NNs) is essential. To do so, HypeRL uses two additional NNs, often called hypernetworks, to learn the weights and biases of the value function and the policy NNs. In this way, HypeRL effectively embeds the parametric information into the value function and policy NNs. We validate the proposed approach on two PDE-constrained optimal control benchmarks, namely a 1D Kuramoto-Sivashinsky equation with in-domain control and on a 2D Navier-Stokes equations with boundary control, by showing that the knowledge of the PDE parameters and how this information is encoded, i.e., via a hypernetwork, is an essential ingredient for learning parameter-dependent control policies that can generalize effectively to unseen scenarios and for improving the sample efficiency of such policies.
@@ 1 Introduction
In this respect, reinforcement learning (RL) [3] is emerging as a new paradigm to address the solution of PDE-constrained OC problems and has been shown to outperform other OC strategies when the system’s states are high-dimensional, noisy, or only partial measurements are available. RL avoids to solve the HJB or the adjoint equations explicitly, which would be untractable for extremely complex problems. Unlike the aforementioned OC approaches, RL aims to solve control problems by learning an OC law (the policy), while interacting with the dynamical system (the environment). RL assumes no prior knowledge of the system, thus yielding broadly applicable control approaches. Deep reinforcement learning (DRL) is the extension of RL using deep neural networks (NNs) to represent value functions and policies [4, 5, 6]. DRL has shown outstanding capabilities in complex control problems such as games [7, 8, 9, 10, 11, 12], simulated and real-world robotics [13, 14, 15, 16, 17, 18, 19], and recently PDEs, with particular emphasis on fluid dynamics [20, 21, 22, 23, 24, 25, 26, 27].
@@
Figure 1: HypeRL for parametric PDE-constrained OC. We rely on a hypernetwork $h(\bm{\mu};\bm{\theta}_{h_{\pi}})$ to learn, from the PDE parameters $\bm{\mu}$, the weights and biases of the policy (and value function) neural network.
@@
Hypernetworks [36] are a class of NNs that provide the parameters, i.e., the weights and biases, of other NNs, often referred to as main or primary networks. Hypernetworks have shown promising results in a variety of deep learning problems, including continual learning, causal inference, transfer learning, weight pruning, uncertainty quantification, zero-shot learning, natural language processing, and recently DRL [37]. Indeed, hypernetworks are capable of enhancing the flexibility, expressivity, and performance of deep learning-based architecture, opening new doors for the development of novel and more advanced architectures. Hypernetworks in DRL were first used in [38] to learn the parameters of the value function or of the policy NNs. Enhancing DRL with hypernetworks has been done in the context of meta RL, zero-shot RL, and continual RL [39, 40, 41] for improving the performance of RL agents. However, to the best of our knowledge, no one has tackled the problem of controlling parametric PDEs with hypernetworks and DRL so far.
In this paper, we propose a novel parameter-informed DRL framework, namely HypeRL, for the efficient solution of parameter-dependent PDE-constrained OC problems by addressing the two aforementioned limitations, namely sample efficiency and generalization. In particular, with reference to Figure 1, we exploit the knowledge of PDE parameters $\bm{\mu}$ to learn a parametrization of the control policy (and value function), dependent on the parameters of the PDE, by means of a hypernetwork $h(\bm{\mu};\bm{\theta}_{h_{\pi}})$ taking as input the PDE parameters $\bm{\mu}$ and providing as output the weights and biases $\bm{\theta}_{\pi}$ of the policy $\pi(\bm{y};\bm{\theta}_{\pi})$ (and value function). In contrast with the simple and widely-used concatenation of the parameters $\mu$ to the PDE state $\bm{y}$, this parametrization allows for learning OC strategies with less data and that can better adapt to unseen instances of the PDE parameters, within and without the training range, i.e., interpolation and extrapolation, respectively. Additionally, we show that the knowledge of the PDE parameters is crucial for learning OC strategies, and that parameter-informed DRL outperforms parameter-unaware DRL.
@@
The paper is organized as follows: in Section 2, we introduce the building blocks of our framework, namely OC, RL, and hypernetworks. In Section 3, we describe our parameter-informed HypeRL framework combining DRL with hypernetworks, and in Section 4 we show the results and discuss the findings. Some conclusions are finally reported in Section 5.
@@ 2.1 Optimal Control Problems
$\begin{split}\min_{\bm{y}\in\mathcal{Y},~{}\bm{u}\in\mathcal{U}_{\text{ad}}}\ \ \ &{J}(\bm{y},\bm{u})\\ \text{subject to ~{}~{}~{}~{}}&\frac{\partial\bm{y}}{\partial t}(t)=\bm{F}(\bm{y}(t),\bm{u}(t);\bm{\mu})\quad\text{for}\;t\in[t_{0},t_{f}]\,,\;\text{with}\;\bm{y}(t_{0})=\bm{y}_{0}\,,\\ \end{split}$ (1)
@@
$\begin{split}L:\mathbb{R}^{|\bm{y}|}\times U_{\text{ad}}\times[t_{0},t_{f}]\rightarrow\mathbb{R},\quad(\bm{x},\bm{z},\tau)\mapsto\frac{1}{2}||\bm{x}-\bm{y}_{\text{ref}}(\tau)||^{2}+\frac{\alpha}{2}||\bm{z}-\bm{u}_{\text{ref}}(\tau)||^{2}\,,\\ \end{split}$ (3)
@@
Note that the state equation is satisfied over the time interval $[t,t_{f}]$ with the state starting with $\bm{y}(t)$. To note, such a value function is the solution to the Hamilton-Jacobi-Bellman (HJB) [43, 44] equation:
$-\frac{\partial V(\bm{y},t)}{\partial t}=\min_{\bm{u}(t)\in U_{\text{ad}}}\left\{L(\bm{y},\bm{u}(t),t)+\bm{F}(\bm{y},\bm{u}(t);\bm{\mu})^{T}\frac{\partial V(\bm{y},t)}{\partial\bm{y}}\right\},\quad(\bm{y},t)\in\mathbb{R}^{|\bm{y}|}\times[t_{0},t_{f}]\,,$ (5)
@@
The HJB equation is typically a high-dimensional PDE as $|\bm{y}|$ is often large. Moreover, due to the dependency of $\bm{F}(\cdot)$ on the (many) parameters $\bm{\mu}$, the HJB equation has to be solved for each sampling location of the parameters, because each sample defines an individual optimal control problem for the specific $\bm{\mu}$. Therefore, (5) is not easily tractable and may be computationally prohibitive for high-dimensional OC problems constrained by PDEs. This trait is commonly known as the curse of dimensionality [3]. We also note that an alternative approach to solving the constrained OC problem (1) uses the Karush-Kuhn-Tucker (KKT) optimality conditions [45, 46] via the introduction of a Lagrange multiplier. However, also in this case the iterative nature of the optimization methods (like, e.g., gradient-based, Newton, quasi-Newton, sequential quadratic programming) makes the numerical solution of each PDE-constrained optimization problem usually very hard. In the case of multiple OC problems, for different parameter values, the overall computational cost would be therefore prohibitive.
Remark 1
To solve (5) one possibility is to rely on traditional numerical methods, e.g., the finite element method. In particular, in the latter case, the semi-discretized form of the HJB equation reads as follows. Let $V(\bm{y},t)$ be linearly approximated by a set of basis functions on $\mathbb{R}^{|\bm{y}|}$, i.e., $V(\bm{y},t)\approx\sum_{j=1}^{|\bm{v}|}v_{j}(t)\phi_{j}(\bm{y})$, in which $\phi_{j}$ is the $j$-th basis function, and $\bm{v}(t)=\{v_{1}(t),\cdots,v_{|\bm{v}|}(t)\}^{T}$ collects the expansion coefficients. Using the same test functions as the basis functions, the Galerkin scheme gives a semi-discretized form of the HJB equation as follows: for $t\in[t_{0},t_{f}]$ and $i=1,\cdots,|\bm{v}|$,
[…]
which has to be solved for $\bm{v}(t)$. We highlight that the solution of (6) entails several issues which need to be addressed. Indeed, the value function is expressed in terms of basis functions depending on the state solution; thus requiring $\bm{y}$ to be known at each time instance. Moreover, handling the right-hand side of (6), that is the minimum in the variational formulation, requires special treatment as, for example, solving local optimization problems at quadrature points as in [47]. As a result, the HJB equation is not easily tractable and may become extremely computationally expensive for high-dimensional and large-time horizon control problems.
@@ 2.2 From Dynamic Programming to Reinforcement Learning
Solving the HJB equation in (5) requires reformulating the problem in a time-discrete framework. Algorithms solving the fully-discretized HJB are typically referred to as dynamic programming (DP) approaches [3, 48]. DP methods can handle all kinds of hybrid systems, even with non-differentiable dynamics, and stochastic OC problems [49]. Examples of DP algorithms are policy iteration and value iteration [3]. DP utilizes the Markov Decision Process (MDP) as underlying mathematical framework in order to account for stochastic systems’ dynamics and tackle a broader class of stochastic OC problems [48].
A MDP is a tuple $\langle Y,U_{\text{ad}},\bm{T},R\rangle$ where $Y\subset\mathbb{R}^{|\bm{y}|}$ is the set of observable states, $U_{\text{ad}}\subset\mathbb{R}^{|\bm{u}|}$ is the set of admissible actions, $\bm{T}:Y\times Y\times U_{\text{ad}}\longrightarrow[0,1]^{|\bm{y}|}$ such that $(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})\longmapsto\bm{T}(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})$ is the transition function, and $R:Y\times U_{\text{ad}}\longrightarrow\mathbb{R}$ such that $(\bm{y}_{k},\bm{u}_{k})\longmapsto R(\bm{y}_{k},\bm{u}_{k})$ denotes the reward function. The transition function $(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})\longmapsto\bm{T}(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})$ describes the probability of reaching state $\bm{y}_{k+1}$ from state $\bm{y}_{k}$ while taking action $\bm{u}_{k}$,
$p(\text{Y}_{k+1}=\bm{y}_{k+1}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k})\,,$ (7)
@@
$\sum_{\bm{y}_{k+1}\in Y}p(\text{Y}_{k+1}=\bm{y}_{k+1}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k})=1,\qquad\forall\ \ \bm{y}_{k}\in Y,\bm{u}_{k}\in U_{\text{ad}}.$ (8)
It is worth mentioning that a deterministic transition function $\bm{y}_{k+1}=\bm{T}(\bm{y}_{k},\bm{u}_{k})$ is a special case arising when $p(\bm{y}_{k+1}|\bm{y}_{k},\bm{u}_{k})=1$ and that the reward function is the fully-discretized counterpart of the running cost $L(\cdot)$, but with opposite sign. Therefore, the value function in DP problems is typically written as a maximization problem over the possible controls rather than a minimization one:
@@
The key idea of DP is to estimate the value function $V(\cdot)$ using the perfectly-known environment dynamics $R(\cdot)$ and $\bm{T}(\cdot)$ and then use it to structure the search for good control strategies. Despite their success, DP algorithms (i) require perfect knowledge of $R(\cdot)$ and $\bm{T}(\cdot)$ – despite they are, in many scenarios, often not known exactly or extremely expensive to compute, especially when the dimensionality of the state is very high – and (ii) are unpractical to use in problems with large number of states due to the need of solving the HJB for all possible states. These two drawbacks drastically limit the application of DP algorithms to complex and large-scale problems such as those arising in the OC of parametric PDEs.
[…]
$G_{k}=r_{k}+\gamma r_{k+1}+\gamma^{2}r_{k+2}+\cdots=\sum_{j=0}^{H}\gamma^{j}r_{k+j}\,,$ (10)
@@
$Q(\bm{y}_{k},\bm{u}_{k})=\mathbb{E}_{\pi^{*}}[G_{k}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k}].$ (12)

### v2

# HypeRL: Hypernetwork-Based Reinforcement Learning for Control of Parametrized Dynamical Systems

@@ Abstract.
In this work, we devise a new, general-purpose reinforcement learning strategy for the optimal control of parametric dynamical systems. Such problems frequently arise in applied sciences and engineering and entail a significant complexity when control and/or state variables are distributed in high-dimensional space or depend on varying parameters. Traditional numerical methods, relying on either iterative minimization algorithms – exploiting, e.g., the solution of the adjoint problem – or dynamic programming – also involving the solution of the Hamilton-Jacobi-Bellman (HJB) equation – while reliable, often become computationally infeasible. Indeed, in either way, the optimal control problem has to be solved for each instance of the parameters, and this is out of reach when dealing with high-dimensional time-dependent and parametric dynamical systems. In this paper, we propose HypeRL, a deep reinforcement learning (DRL) framework to overcome the limitations shown by traditional methods. HypeRL aims at approximating the optimal control policy directly, bypassing the need to numerically solve either the HJB equation explicitly for all possible states and parameters, or an adjoint problem within an iterative optimization loop for each parameter instance. Specifically, we employ an actor-critic DRL approach to learn an optimal feedback control strategy that can generalize across the range of variation of the parameters. To effectively learn such optimal control laws for different instances of the parameters, encoding the parameter information into the DRL policy and value function neural networks (NNs) is essential. To do so, HypeRL uses two additional NNs, often called hypernetworks, to learn the weights and biases of the value function and the policy NNs. In this way, HypeRL effectively embeds the parametric information into the value function and policy NNs. We validate the proposed approach on two parametric control problems, namely (i) a 1D parametric Kuramoto-Sivashinsky equation with in-domain control, and (ii) a navigation problem of particle dynamics in a parametric 2D gyre flow. We show that the knowledge of physical and task-dependent information and the encoding of this information via a hypernetwork, are essential ingredients for learning parameter-dependent control policies that can generalize effectively to unseen scenarios, and for improving the sample efficiency of such policies.
@@ 1 Introduction
In this respect, reinforcement learning (RL) [54] is emerging as a new paradigm to address the solution of PDE-constrained OC problems and has been shown to outperform other OC strategies when the system’s states are high-dimensional, noisy, or only partial measurements are available. RL avoids to solve the HJB or the adjoint equations explicitly, which would be untractable for extremely complex problems. Unlike the aforementioned OC approaches, RL aims to solve control problems by learning an OC law (the policy), while interacting with the dynamical system (the environment). RL assumes no prior knowledge of the system, thus yielding broadly applicable control approaches. Deep reinforcement learning (DRL) is the extension of RL using deep neural networks (NNs) to represent value functions and policies [1, 36, 17]. DRL has shown outstanding capabilities in complex control problems such as games [43, 66, 52, 33, 42, 58], simulated and real-world robotics [38, 29, 45, 67, 20, 68, 7], and recently PDEs, with particular emphasis on fluid dynamics [10, 4, 11, 15, 46, 65, 44, 69, 8].
@@
Figure 1: HypeRL for parametric PDE-constrained OC. We rely on a hypernetwork $h_{\pi}(\bm{z}_{k};\bm{\theta}_{h_{\pi}})$ to learn the weights and biases of the policy neural network.
@@
Hypernetworks [22] are a class of NNs that provide the parameters, i.e., the weights and biases, of other NNs, often referred to as main or primary networks. Hypernetworks have shown promising results in a variety of deep learning problems, including continual learning, causal inference, transfer learning, weight pruning, uncertainty quantification, zero-shot learning, natural language processing, and recently DRL [12]. Indeed, hypernetworks are capable of enhancing the flexibility, expressivity, and performance of deep learning-based architecture, opening new doors for the development of novel and more advanced architectures. Hypernetworks in DRL were first used in [49] to learn the parameters of the value function or of the policy NNs. Enhancing DRL with hypernetworks has been done in the context of meta RL, zero-shot RL, and continual RL [3, 47, 25] for improving the performance of RL agents. However, to the best of our knowledge, no one has tackled the problem of controlling parametric dynamical systems with hypernetworks and DRL so far.
In this paper, we propose HypeRL, a novel DRL algorithm for control parameter-dependent PDE-constrained OC problems tailored to address the two aforementioned limitations of traditional DRL approaches, namely sample efficiency and generalization. In particular, with reference to Figure 1, as per standard RL settings, we discretize in time and space the environment dynamics described by:
[…]
where $T(\bm{y}_{k},\bm{u}_{k};\bm{\mu})$ and $R(\bm{y}_{k},\bm{u}_{k};\bm{\mu})$ are the transition and reward functions, $\bm{y}_{k+1},\bm{y}_{k}$ indicate the state at timestep $k+1$ and $k$, respectively, $\bm{u}_{k}$ the control input, $r_{k+1}$ is the reward, and $\bm{\mu}$ the parametric dependency of the environment, which might include physical or task-dependent information. In addition, we define the agent’s state at timestep $k$ as $\bm{z}_{k}=[\bm{y}_{k},\bm{\mu}]$. We exploit the knowledge of parameters $\bm{\mu}$ to learn a parametrization of the control policy $\pi(\bm{z}_{k};\bm{\theta}_{\pi})$ and value function $Q(\bm{z}_{k},\bm{u}_{k};\bm{\theta}_{Q})$ by means of two hypernetworks $h_{\pi}(\bm{z}_{k};\bm{\theta}_{h_{\pi}})$
[…]
and $h_{Q}(\bm{z}_{k};\bm{\theta}_{h_{Q}})$
[…]
where $\bm{\theta}_{h_{\pi}}$ and $\bm{\theta}_{h_{Q}}$ indicate the hypernetwork learnable parameters, $\bm{\theta}_{\pi}$ the policy learnable parameters, and $\bm{\theta}_{Q}$ the value function ones. We test our hypernetwork-based method on two challenging parametric control problems, namely (i) the stabilization of a parametric Kuramoto-Sivashinsky equation to arbitrary references and (ii) the control of a particle to arbitrary target locations in a parametric gyre flow. Throughout our experiments, we show that the encoding via hypernetworks of the physical and task-dependent information enables effective coordination and enhances generalization capabilities compared to traditional and state-of-the-art DRL algorithms.
@@
The paper is organized as follows: in Section 2, we introduce the building blocks of our framework, namely OC, RL, and hypernetworks. In Section 3, we describe our parameter-informed HypeRL framework combining DRL with hypernetworks, and in Section 4 we show the results and discuss the findings. The conclusions are finally reported in Section 5.
@@ 2.1 Optimal Control Problems
[…]
@@
[…]
@@
Note that the state equation is satisfied over the time interval $t,t_{f}]$ with the state starting with $\bm{y}(t)$. To note, such a value function is the solution to the Hamilton-Jacobi-Bellman (HJB) [28, 16] equation:
$\frac{\partial V(\bm{y},t)}{\partial t}=\min_{\bm{u}(t)\in U_{\text{ad}}}\left\{L(\bm{y},\bm{u}(t),t)+\bm{F}(\bm{y},\bm{u}(t);\bm{\mu})^{T}\frac{\partial V(\bm{y},t)}{\partial\bm{y}}\right\},\quad(\bm{y},t)\in\mathbb{R}^{|\bm{y}|}\times[t_{0},t_{f}]\,$, (5)
@@
The HJB equation is typically a high-dimensional PDE as $|\bm{y}|$ is often large. Moreover, due to the dependency of $\bm{F}(\cdot)$ on the (many) parameters $\bm{\mu}$, the HJB equation has to be solved for each sampling location of the parameters, because each sample defines an individual optimal control problem for the specific $\bm{\mu}$. Therefore, (5) is not easily tractable and may be computationally prohibitive for high-dimensional OC problems constrained by PDEs. This trait is commonly known as the curse of dimensionality [54]. We also note that an alternative approach to solving the constrained OC problem (1) uses the Karush-Kuhn-Tucker optimality conditions [27, 32] via the introduction of a Lagrange multiplier. However, also in this case the iterative nature of the optimization methods (like, e.g., gradient-based, Newton, quasi-Newton, sequential quadratic programming) makes the numerical solution of each PDE-constrained optimization problem usually very hard. In the case of multiple OC problems, for different parameter values, the overall computational cost would be therefore prohibitive.
@@ 2.2 From Dynamic Programming to Reinforcement Learning
Solving the HJB equation in (5) requires reformulating the problem in a time-discrete framework. Algorithms solving the fully-discretized HJB are typically referred to as dynamic programming (DP) approaches [54, 5, 6]. DP methods can handle all kinds of hybrid systems, even with non-differentiable dynamics, and stochastic OC problems [14]. Examples of DP algorithms are policy iteration and value iteration [54]. DP utilizes the Markov Decision Process (MDP) as underlying mathematical framework in order to account for stochastic systems’ dynamics and tackle a broader class of stochastic OC problems [5].
A MDP is a tuple $\langle Y,U_{\text{ad}},T,R\rangle$ where $Y\subset\mathbb{R}^{|\bm{y}|}$ is the set of observable states, $U_{\text{ad}}\subset\mathbb{R}^{|\bm{u}|}$ is the set of admissible actions, $T:Y\times Y\times U_{\text{ad}}\longrightarrow[0,1]^{|\bm{y}|}$ such that $(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})\longmapsto T(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})$ is the transition function, and $R:Y\times U_{\text{ad}}\longrightarrow\mathbb{R}$ such that $(\bm{y}_{k},\bm{u}_{k})\longmapsto R(\bm{y}_{k},\bm{u}_{k})$ denotes the reward function. The transition function $(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})\longmapsto T(\bm{y}_{k+1},\bm{y}_{k},\bm{u}_{k})$ describes the probability of reaching state $\bm{y}_{k+1}$ from state $\bm{y}_{k}$ while taking action $\bm{u}_{k}$,
$p(\text{Y}_{k+1}=\bm{y}_{k+1}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k})\,$,
@@
$\sum_{\bm{y}_{k+1}\in Y}p(\text{Y}_{k+1}=\bm{y}_{k+1}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k})=1,\qquad\forall\ \ \bm{y}_{k}\in Y,\bm{u}_{k}\in U_{\text{ad}}.$
It is worth mentioning that a deterministic transition function $\bm{y}_{k+1}=T(\bm{y}_{k},\bm{u}_{k})$ is a special case arising when $p(\bm{y}_{k+1}|\bm{y}_{k},\bm{u}_{k})=1$ and that the reward function is the fully-discretized counterpart of the running cost $L(\cdot)$, but with opposite sign. Therefore, the value function in DP problems is typically written as a maximization problem over the possible controls rather than a minimization one:
@@
The key idea of DP is to estimate the value function $V(\cdot)$ using the perfectly-known environment dynamics $R(\cdot)$ and $T(\cdot)$ and then use it to structure the search for good control strategies. Despite their success, DP algorithms (i) require perfect knowledge of $R(\cdot)$ and $T(\cdot)$ – despite they are, in many scenarios, often not known exactly or extremely expensive to compute, especially when the dimensionality of the state is very high – and (ii) are unpractical to use in problems with large number of states due to the need of solving the HJB for all possible states. These two drawbacks drastically limit the application of DP algorithms to complex and large-scale problems such as those arising in the OC of parametric PDEs.
Reinforcement learning (RL) [54] is a promising machine learning approach to solve sequential decision-making problems through a trial-and-error process. Unlike DP, RL does not try to solve the HJB for all possible states; rather, it aims at deriving OC laws from (i) measurements of the system – often referred to as observations, and (ii) reward samples, without direct knowledge of $T(\cdot)$ and $R(\cdot)$ [54, 35]. In RL, we can identify two main entities: the agent and the environment (see Figure 1). The agent aims to find the best strategy to solve a given task by interacting with an unknown environment. Similarly to DP, the optimality of the strategy learned by the agent is defined by a task-dependent reward function. In particular, we can use RL to solve OC problems, such as the one in Equation (1), by learning an OC law from data without the need to explicitly solve the HJB equation (either in continuous or in discrete settings) for all possible states. Similarly to DP, we can rely on MDPs to mathematically formulate the RL problem. The goal of any RL algorithm is to find the optimal policy (control law) maximizing the expected cumulative return $G_{k}$:
$G_{k+1}=r_{k+1}+\gamma r_{k+2}+\gamma^{2}r_{k+3}+\cdots=\sum_{j=1}^{H}\gamma^{j}r_{k+j}\,$, (7)
@@
$Q(\bm{y}_{k},\bm{u}_{k})=\mathbb{E}_{\pi^{*}}[G_{k+1}|\text{Y}_{k}=\bm{y}_{k},\text{U}_{k}=\bm{u}_{k}].$ (9)